OpenStudy (anonymous):

((x^(2/5)/9y(-16))^1/2

jhonyy9 (jhonyy9):

-16 is exponent ?

jhonyy9 (jhonyy9):

so x will be on exponent 2/5 * 1/2 = 2/10 = 1/5

jhonyy9 (jhonyy9):

and y on exponent -8